OpenStudy (anonymous):

simple exponential equations problem! determine the exact value of x 3^(2x)= 5(3^x)+ 36

OpenStudy (dumbcow):

treat it like a quadratic u = 3^x _>u^2-5u-36 = 0 Factor and solve for u Then replace u with 3^x and solve for x

OpenStudy (anonymous):

ok so i factored and go (u-9)(u+4).. i dont really understand what to do next!

OpenStudy (dumbcow):

good solve each part separately u-9 =0 u+4=0

OpenStudy (anonymous):

ok, so u=9 and -4

OpenStudy (dumbcow):

which means 3^x = 9 and 3^x = -4 Find x

OpenStudy (anonymous):

but the left side of the equation is 3^(2x), not 3^x...

OpenStudy (dumbcow):

correct but took care of that part when we made it a quadratic 3^2x became u^2 dont worry its the same x

OpenStudy (anonymous):

ok so for 3^x=9, x = 2 but i dont know how to do it for the other one

OpenStudy (dumbcow):

good actually the one doesnt exist 3^x > 0

OpenStudy (dumbcow):

graph it to see for yourself

OpenStudy (dumbcow):

there is only one answer of 2

OpenStudy (anonymous):

ok awesome! thank you sososo much!

OpenStudy (dumbcow):

welcome